OpenStudy (anonymous):

I really need help with systems of equationss:) 4x-y=8 -8x+2y=-5

OpenStudy (anonymous):

4x+y=2

OpenStudy (anonymous):

What do you want? Want to find X and Y?

OpenStudy (cwrw238):

you can solve these by elimination method eliminate the term in y by multiplying the first equation by 2: this equation becomes 8x - 2y = 16 now add this to the second equation -8x + 2y = -5 and you get 0 = 11 which of course is rubbish

OpenStudy (anonymous):

yes, exactly!

OpenStudy (cwrw238):

so what can we conclude from this ?

OpenStudy (anonymous):

that this equation has no solution? I dont know :(

OpenStudy (cwrw238):

yes - no solutions

OpenStudy (cwrw238):

sometimes you can see this by observation - i missed it this time

OpenStudy (anonymous):

Gave my try and got... yeah - this does not seem valid.

OpenStudy (anonymous):

thank you so much for your help!!!!

OpenStudy (cwrw238):

yw